OpenStudy (idealist10):

Find the general solution of (2x+1)y"-2y'-(2x+3)y=(2x+1)^2, given that y1=e^-x satisfies the complementary equation.

OpenStudy (idealist10):

y=ue^-x y'=u'e^(-x)-ue^(-x) y"=u"e^(-x)-2u'e^(-x)+ue^(-x) Subbing: (2x+1)y"-2y'-(2x+3)y=(2x+1)(u"e^(-x)-2u'e^(-x)+ue^(-x))-2(u'e^(-x)-ue^(-x))-(2x+3)ue^(-x)=u"2xe^(-x)+u"e^(-x)-4u'xe^(-x)-4u'e^(-x)=e^(-x)(u"2x+u"-4u'x-4u') So, e^(-x)(u"2x+u"-4u'x-4u')=(2x+1)^2 u"2x+u"-4u'x-4u'=e^(x)(2x+1)^2 Am I right up to this point? @wio

OpenStudy (anonymous):

Hard to read

OpenStudy (anonymous):

I'm not sure, actually. I haven't done this in a while

OpenStudy (idealist10):

@ganeshie8

OpenStudy (anonymous):

I think you're correct up to that step

OpenStudy (idealist10):

Okay, so what should I do next then?

OpenStudy (anonymous):

What other methods do you know?

OpenStudy (anonymous):

They don't teach you want to do when you have variable coefficients?

OpenStudy (idealist10):

No, it's because at the end, I got some weird results, that's why. If you said up to that point I'm right, then this is what I got next from above. u"(2x+1)-4u'(x+1)=e^(x)(2x+1)^2 \[u''-\frac{ 4u'(x+1) }{ 2x+1 }=e ^{x}(2x+1)\]

OpenStudy (idealist10):

Then z=u', subbing: \[z'-\frac{ 4(x+1) }{ 2x+1 }z=e ^{x}(2x+1)\]

OpenStudy (idealist10):

Up to this point above, I tried to integrate the middle term because I applied the method of integrating factor in order to find z at the end and integrate u in order to find y.

OpenStudy (idealist10):

Any other method to solve for z without doing the integrating factor method?

OpenStudy (inkyvoyd):

Seems like you should just bite the bullet and do your partial fractions